OpenStudy (anonymous):

dy/dx of xy^2+4y-5x^2 (implicit differentation) plz explain how to do this problem step by step...plz and ty..

OpenStudy (anonymous):

We differentiate with respect to 'x' if for one term there is a 'y' we differentiate it and multiply it by dy/dx or y'. So for your question: Using implicit differentiation: Differentiate xy^2 using the chain rule: y^2+2xy*(dy/dx). Do you understand this?

OpenStudy (anonymous):

yes but i would appreciate if u could work out the porblem so i can use it as an example on similar problems like this that could be on my exam tmmrw and i dnt fail that exam..

OpenStudy (kainui):

\[xy^2+4y-5x^2=0\] Is this really your equation?

OpenStudy (anonymous):

Differentiate xy^2: Using chain rule: differentiate with respect to x and treat the y term as a constant (y^2) then differentiate with respect to y and treat the x term as a constant (2xy*dy/dx). Ans.: y^2+2xy*dy/dx. Differentiate 4y: Because we are asked to differentiate with respect to x if we differentiate with respect to y we multiply our answer by dy/dx always. Ans.: 4*dy/dx. Differentiate 5x^2 = 10x. Complete answer: y^2 + 2xy*dy/dx+4*dy/dx -10x.

OpenStudy (kainui):

I always look at "y" as though it's derivative is "dy/dx". This helps when looking at the chain rule when taking the derivative of the outside times the derivative of the inside. For example, the derivative of: 3y^5 is: 15y^5*(dy/dx) Does that explain it? Then what you do is solve for (dy/dx) after you've differentiated with algebra similar to how you would normally solve for something like x.

OpenStudy (kainui):

its*

OpenStudy (anonymous):

derivative of 3y^5 is 15y^4 dy/dx

OpenStudy (kainui):

lol my bad, but yes you get the point?

OpenStudy (anonymous):

\[(uv)'=uv'+u'v\]

OpenStudy (anonymous):

That is the product rule

OpenStudy (anonymous):

yup u are right it is the product rule not the chain rule. @ChmE

OpenStudy (anonymous):

But you were also right in saying that it is also the chain rule. Both are used

OpenStudy (anonymous):

Implicit Differentiation: Hmm, how should I explain it. Basically it is showing the chain rule instead of it just staying hidden. The derivative of y^2 is 2y. I think we all would agree on that but what don't we see? It is really y^2 * y' and since y' = 1 we just leave that out. But if wanted to take the derivative of (2x+1)^2 it would be 2(2x+1) * 2 because the derivative of y (in this case y=2x + 1) is equal to 2, so it is not hidden here. All implicit differentiation really is, is showing the chain rule as we differentiate. y' is the same as dy/dx. dy/dx means that we are taking the derivative of y with respect to x. So after every derivative of y we have to include dy/dx after it. The only explanation for why we don't do dx/dy after the x is because we are only trying to solve for dy/dx. If anybody has a better explanation for this part plz feel free to add. After we differentiate we need to solve for dy/dx. This is just all algebra. Ex) 3x^2 + 2y^3 = 0 6x + 6y^2 dy/dx = 0 Now we have to solve for dy/dx though 6y^2 dy/dx = -6x dy/dx = -6x/6y^2 dy/dx = -x/y^2

OpenStudy (anonymous):

\[x \color{blue}{y^2}+4 \color{blue}{y}-5x^2=0\]\[2x \color{blue}{y} \color{red}{\frac{dy}{dx}}+\color{blue}{y^2}+4 \color{red}{\frac{dy}{dx}}-10x=0\]\[\color{red}{\frac{dy}{dx}}(2x \color{blue}{y}+4)=10x-\color{blue}{y^2}\]\[\color{red}{\frac{dy}{dx}}=\frac{10x-\color{blue}{y^2}}{2x \color{blue}{y}+4}\]

zepdrix (zepdrix):

lolol nice :)

OpenStudy (anonymous):

Oh it's on now. Colors are the best!!

OpenStudy (agent0smith):

"The only explanation for why we don't do dx/dy after the x is because we are only trying to solve for dy/dx." Wouldn't it make more sense to say, we don't use dx/dy because we're differentiating the equation with respect to x? dx/dy would mean we're differentiating the equation with respect to y.

OpenStudy (anonymous):

Nice I like that